### v2

# Transplant Then Regenerate: A New Paradigm for Text Data Augmentation

Abstract
Data augmentation is a critical technique in deep learning. Traditional methods like Back-translation typically focus on lexical-level rephrasing, which primarily produces variations with the same semantics. While large language models (LLMs) have enhanced text augmentation by their “knowledge emergence” capability, controlling the style and structure of these outputs remains challenging and requires meticulous prompt engineering. In this paper, we propose LMTransplant, a novel text augmentation paradigm leveraging LLMs. The core idea of LMTransplant is transplant-then-regenerate: incorporating seed text into a context expanded by LLM, and asking the LLM to regenerate a variant based on the expanded context. This strategy allows the model to create more diverse and creative content-level variants by fully leveraging the knowledge embedded in LLMs, while preserving the core attributes of the original text. We evaluate LMTransplant across various text-related tasks, demonstrating its superior performance over existing text augmentation methods. Moreover, LMTransplant demonstrates exceptional scalability as the size of augmented data grows.
Transplant Then Regenerate: A New Paradigm for Text Data Augmentation
Guangzhan Wang   Hongyu Zhang   Beijun Shen   Xiaodong Guthanks: Correspondence: xiaodong.gu@sjtu.edu.cn Shanghai Jiao Tong UniversityChongqing University {wangguangzhan, bjshen, xiaodong.gu}@sjtu.edu.cn hyzhang@cqu.edu.cn
1 Introduction
Data augmentation is a critical technique in deep learning khosla2020enhancing; DA_DL_2021; DA_DL_2022. Deep learning models are data-hungry and often suffer from limited datasets. Data augmentation generates additional training samples through transforming or rephrasing the existing dataset. This process increases data diversity, reduces the risk of overfitting, and enhances the models’ generalization ability.
Data augmentation has been extensively studied in NLP tasks feng-etal-2021-survey; 2023-survey; chen-etal-2023-survey; DA-2023-survey. One simple prior approach wei-zou-2019-eda enhances text data through word-level transformations, such as random insertion and deletion. While easy to implement, it often generates low-quality samples that severely disrupt the semantic coherence of the generated text. Later, Back-translation 2015Improving has been commonly used for sentence level rephrasing. Specifically, a translation model first translates the original text into a different language and then translates it back into the original language. Although the augmented text is semantically coherent, it often exhibits high similarity to the original text, leading to poor data diversity 2023-survey; understanding.
Figure 1: Illustration of LMTransplant.
Recent advancements in LLMs have spurred significant interest in LLM-based data augmentation methods whitehouse-etal-2023-llm; ding-etal-2024-data; zhou2024survey; qiao2024autoact. A key characteristic of LLMs is their “knowledge emergence” capability, which stems from two factors: (1) the extensive prior knowledge acquired during pre-training, and (2) their robust language understanding and instruction-following abilities CoDa; ABEX. These strengths allow LLMs to generate desired outputs directly through demonstrations or natural language instructions without requiring additional training. A notable example is AugGPT dai2023auggpt, which instructs ChatGPT to rephrase text for improving text classification performance. However, this rephrasing-based method only generates variants with similar semantics and underutilizes the rich knowledge embedded in LLMs, limiting the diversity and creativity of the generated data. As a result, when augmented data volume reaches a certain threshold, further increasing may not yield performance improvements zhou2024survey. Although contextual augmentation methods can enhance content diversity, such as GPT3Mix yoo-etal-2021-gpt3mix-leveraging, which leverages the powerful few-shot learning capabilities of LLMs to generate mixed augmented text, it is highly sensitive to example quality and input order, potentially introducing uncontrolled biases that are not aligned with the original data distribution.
To address these limitations and better leverage the potential of LLMs, we propose LMTransplant, a novel text data augmentation paradigm. The core idea of LMTransplant is transplant-then-regenerate (TTR), namely, embedding the original text into an expanded context generated by LLMs and then instructing LLMs to regenerate a variant based on this enriched context. Specifically, LMTransplant employs bidirectional text continuation—backward (right-to-left) and forward (left-to-right)—to create the preceding and subsequent context of the original text. This original text is then masked within its expanded context, and LLMs are prompted to generate a replacement that introduces novel content diversity beyond rephrasing, while still preserving core attributes of the original text. Therefore, LMTransplant enhances both the diversity and creativity of the generated text, while maintaining alignment with the original data distribution.
We apply LMTransplant to various deep learning tasks, including text classification, question answering and named entity recognition (NER), and compare its performance with existing data augmentation methods. Experimental results demonstrate that LMTransplant can generate higher-quality augmented data. Training models with data augmented by our approach yields significant performance improvements across all tasks. Compared to non-augmentation, LMTransplant achieves accuracy gains of 28.16%, 19.96%, 7.68%, 23.66%, and 10.25% on the SST-2, TREC, SNIPS, MLQA, and CoNLL-2003 datasets, respectively.
In summary, our contributions are as follows:
  • We propose a novel transplant-based paradigm for text data augmentation. Unlike existing methods which primarily focus on rephrasing, LMTransplant crafts content-level text variants, thereby crafting higher-quality augmented texts.
  • We present a novel transplant and regeneration algorithm using bidirectional text continuation and masked text prediction. The algorithm allows for generating core attributes similar yet more diverse and creative text by effectively utilizing knowledge embedded in LLMs. Experiments demonstrate that, LMTransplant achieves significant performance improvements across different tasks.
2 Approach
2.1 Overview
We propose LMTransplant, a novel data augmentation paradigm. The core idea of LMTransplant is transplant-then-regenerate: integrating the original text into contextual scenarios generated by LLMs, and then asking LLMs to regenerate new variants given expanded contexts. This strategy allows the model to create content-level variants while preserving the core attributes of the original text. Figure 1 illustrates the entire augmentation process. For a given text, LMTransplant uses bidirectional text continuation to generate its preceding and subsequent contexts (Section 2.2). Subsequently, LMTransplant masks the original text in the transplanted text and asks the LLM to regenerate the missing parts given the crafted contexts, thereby producing new variants of the original text (Section 2.3). Each step of this process will be elaborated in the following sections.
2.2 Transplant
Given a seed text, we incorporate it into a relevant contextual scenario. Specifically, we treat the seed text as a fragment of a broader contextual passage, and then use an LLM to generate a semantically natural and logically coherent surrounding context. This process can be conceptualized as bidirectional text continuation, which involves two steps: (1) a forward (left-to-right) continuation process that continues writing the subsequent context of the seed text, and (2) a backward (right-to-left) continuation process that reconstructs the preceding context of the seed text.
Figure 2: Illustration of text transplant.
Figure 2 illustrates the bidirectional text continuation process. First, the LLM generates a subsequent text that naturally extends the seed text, combining both to form an expanded passage. Next, the LLM generates a preceding text based on this expanded input. This process ensures that the generated context remains thematically aligned with the original text, while introducing novel information other than rephrasing through the knowledge embedded in LLMs, such as new words and expressions. Leveraging the language understanding and instruction-following capabilities of LLMs, we employ prompt engineering to guide content generation. The prompt template used for transplanting is provided in Appendix LABEL:sec:PromptTemplates.
2.3 Regeneration
The transplant phase generates multiple contextual scenes bearing the original text. We then introduce a regeneration process, where we prompt the LLM to regenerate new text variants that seamlessly integrate into the expanded context. Specifically, we provide the LLM with the crafted preceding and subsequent contexts, along with the original text, and ask it to generate a text that introduces content variation while preserving essential attributes of the original. Therefore, the regenerated text must satisfy the following criteria: (1) Fitting naturally within the surrounding context; (2) Aligning with the original text in terms of theme, length, format, and linguistic style, as mismatches in these aspects between training and testing instances are known to degrade downstream performance rogers2021primer; (3) Introducing novel elements to enrich content variation, avoiding simple rewording or direct replication.
In this step, the preceding and subsequent texts generated in Section 2.2 serve as a bridge, ensuring the newly generated text aligns with the original, such as theme and label-related information. Simultaneously, this process enriches and diversifies the content, making the regenerated text a high-quality augmentation of the original. The prompt design for this step is detailed in Appendix LABEL:sec:PromptTemplates.
Figure 3: Illustration of regeneration.
3 Experiments
We conduct extensive experiments to evaluate the effectiveness of LMTransplant across multiple deep learning tasks by applying data augmentation to various datasets. Following established practices yoo-etal-2021-gpt3mix-leveraging; dai2023auggpt; ubani2023zeroshotdataaug; lee2024llm2llm, we simulate low-resource scenarios by subsampling the training set of each dataset. Specifically, we select a subset of samples as seed data for augmentation, and then generate three augmented samples for each seed. This enables us to rigorously assess the performance of our method in data-scarce scenarios.
3.1 Implementation Details
To evaluate LMTransplant’s robustness and effectiveness across a wider range of LLMs, we employ different LLMs as the base LLM for both transplant and regeneration phases, including DeepSeek-V3, GPT-3.5-Turbo and GPT-4o.
For text classification tasks, we use ModernBERT ModernBERT, a modernized bidirectional encoder-only BERT-style model available in two sizes: ModernBERT-base and ModernBERT-large. The classifier is initialized using pre-trained models from the Huggingface Transformers library Huggingface and optimized with the AdamW optimizer Adam; 2017Decoupled. We set the learning rate to 4e-5 and maintain other hyperparameters consistent with fadaee-etal-2017-data, including 8 training epochs and a batch size of 8. During training, models are saved based on their performance on the development set, with the best-performing parameters retained for final evaluation on the test set.
For the question-answering task, we fine-tune Qwen2.5-1.5B qwen2.5; qwen2 using the AdamW optimizer with a learning rate of 1e-5, a batch size of 8, and 8 training epochs. And for NER task, we fine-tune ModernBERT using the same hyperparameters as in the classification task.
We repeat all experiments 10 times to mitigate the influence of randomness. Additionally, we conduct pairwise two-sample Wilcoxon Signed-Rank tests PairwiseWilcoxonSignedRankTest to compare group medians and assess statistical significant differences. When $p<0.01$, it indicates that there is a statistical significant difference between the two groups.
3.2 Datasets
Task Type Dataset Classes Train Dev Test
Classification SST-2 2 6228 692 1821
TREC 6 5406 546 500
SNIPS 7 13084 700 700
[1pt/2pt] QA MLQA - 1314 437 437
[1pt/2pt] NER CoNLL-2003 9 3234 748 679
Table 1: Statistics of datasets.
We evaluate LMTransplant on five established benchmarks:
  • SST-2 sst2: A widely used sentiment classification dataset of movie reviews, labeled as “positive” or “negative”. We use the version provided by ConditionalBERT [^1] https://github.com/1024er/cbert_aug, which contains 6,228 training samples, 692 development samples, and 1,821 test samples.
  • TREC trec: A question classification dataset annotated with six types: “Abbreviation”, “Entity”, “Description”, “Human”, “Location”, and “Numeric”. Similar to SST-2, we use the version from ConditionalBERT, with 5,406 training samples, 546 development samples, and 500 test samples.
  • SNIPS snips: A text classification dataset annotated with seven human intents: “AddToPlaylist”, “BookRestaurant”, “GetWeather”, “PlayMusic”, “RateBook”, “SearchCreativeWork”, and “SearchScreeningEvent”. We use the version from SlotGated-SLU goo-etal-2018-slot [^2] https://github.com/MiuLab/SlotGated-SLU/tree/master/data/snips, comprising 13,084 training samples, 700 development samples, and 700 test samples.
  • MLQA lewis2019mlqa: A question-answering benchmark with context passages, questions, and answers. We use English samples form Huggingface [^3] https://huggingface.co/datasets/dkoterwa/mlqa_filtered and filter out those exceeding 80 tokens to keep a modest length.
  • CoNLL-2003 CoNLL-2003: A NER dataset contains four entity types: persons, organizations, locations, and miscellaneous names, tagged using the IOB scheme, resulting in nine distinct IOB labels.
Statistics of datasets are summarized in Table 1.
Method SST-2 TREC SNIPS MLQA CoNLL-2003
[…]
Original 0.99±0.01 - 0.94±0.02 - 0.89±0.03 - 0.97±0.01 - 0.97±0.01 -
MoreData 0.99±0.01 - 0.89±0.01 - 0.80±0.02 - 0.93±0.02 - 0.94±0.01 -
[1pt/2pt] EDA 0.31±0.02 0.19±0.02 0.46±0.02 0.23±0.01 0.30±0.01 0.16±0.01 0.40±0.01 0.21±0.01 0.43±0.01 0.23±0.01
BackTrans. 0.69±0.03 0.20±0.01 0.55±0.03 0.13±0.01 0.60±0.02 0.19±0.01 0.53±0.01 0.16±0.06 0.62±0.03 0.12±0.01
GPT3Mix 0.67±0.03 - 0.54±0.02 - 0.50±0.04 - 0.65±0.04 - 0.18±0.02 -
AugGPT 0.54±0.03 0.24±0.01 0.37±0.02 0.19±0.01 0.41±0.02 0.27±0.01 0.38±0.02 0.19±0.01 0.33±0.02 0.22±0.01
LLM2LLM 0.71±0.02 0.23±0.01 0.52±0.03 0.17±0.01 0.54±0.01 0.22±0.01 0.61±0.01 0.17±0.01 0.71±0.02 0.20±0.02
[1pt/2pt]   LMTransplant (ours)
(left, right) 0.88±0.03 0.39±0.01 0.66±0.03 0.30±0.01 0.63±0.02 0.36±0.01 0.72±0.03 0.29±0.01 0.78±0.03 0.27±0.01
(right, left) 0.88±0.03 0.39±0.01 0.64±0.03 0.29±0.01 0.63±0.02 0.36±0.01 0.71±0.02 0.28±0.01 0.75±0.03 0.27±0.02
Unidirectional 0.82±0.03 0.37±0.01 0.54±0.02 0.25±0.01 0.50±0.02 0.31±0.01 0.63±0.02 0.25±0.01 0.66±0.03 0.22±0.01
Table 2: Quality of generated samples by various methods ($p<0.01$). Subscript numbers denote standard deviations. SV == Semantic Variability; MoreData: randomly samples additional data from the original training set as augmented data. Underlines indicate the second-best performance for each metric. We choose DeepSeek-V3 as the base LLM for data augmentation. Results based on GPT-3.5-Turbo and GPT-4o are available in Appendix B.
3.3 Metrics
We evaluate LMTransplant along two dimensions: the quality of the augmented texts (intrinsic evaluation) and their impact on deep learning tasks (extrinsic evaluation). For intrinsic evaluation, we adopt two widely employed metrics to assess the quality of the augmented samples:
Distinct-N measures the lexical diversity DISTINCT. It is defined as the ratio of unique n-grams across all generated texts and their corresponding seeds.
$\text{Distinct-N}=\frac{\#\text{ unique n-grams}}{\#\text{ all n-grams}}$ (1)
A higher value indicates greater diversity. We calculate the average score across 10 experiments as the final result.
Semantic Variability measures how well the generated text extends the semantics of the seed texts. To assess this, we first adopt BERTScore BERTScore to calculate the sentence-level similarity between the generated text and its original, leveraging BERT’s contextual embeddings. We then define semantic variability as:
$\text{Semantic Variability}=1-\text{BERTScore}$ (2)
Thus, the higher the semantic variability, the better the variability that the new text variants are.
For extrinsic evaluation, we adopt task-related metrics. Specifically, we use accuracy (Acc) and macro F1-score (Macro-F1) for classification and NER tasks, and the accuracy of answers (Acc) for QA tasks:
$\text{Acc}=\frac{\#\text{samples answered correctly}}{\#\text{all test samples}}$ (3)
$\text{Macro-F1}=\frac{1}{N}\sum_{i=1}^{N}F1_{i}$ (4)
where $N$ is the total number of classes, and $F1_{i}$ represents the F1-score for the i-th class.
3.4 Baselines
We compare our method with both traditional and LLM-based methods:
Easy Data Augmentation (EDA) wei-zou-2019-eda: A rule-based data augmentation method that applies lexical transformations, including synonym replacement, random insertion, random swap, and random deletion, to the original text.
Back Translation (BackTrans.) 2015Improving: A widely used date augmentation method that translates the original text into another language and then back-translates it into the original language to generate variants. Following ZeroShotDataAug ubani2023zeroshotdataaug, we use googletrans [^4] Google Translate (https://pypi.org/project/googletrans/) as the machine translation model, selecting different intermediate languages for multiple augmentations of the same text.
GPT3Mix yoo-etal-2021-gpt3mix-leveraging: This method randomly selects several examples from the seed samples, embeds them into a prompt template, and then leverages the powerful few-shot learning capabilities of LLMs to generate mixed augmented text influenced by the provided examples.
AugGPT dai2023auggpt: This method utilizes prompts to guide LLMs in rephrasing each sentence from the training samples into multiple semantically similar but linguistically different variants, thereby enhancing text classification performance.
LLM2LLM lee2024llm2llm: An iterative data augmentation strategy that continuously employs LLMs to generate new samples from instances mispredicted by the downstream task model.
Subsequent Sentence: Generate a sentence that can naturally follow the original text, ensuring a smooth transition and logical continuation. New Text: Place the subsequent sentence behind the original <text_type> to form a new text. Preceding Sentence: Create a sentence that can naturally precede the new <text_type>, making the transition smooth and logical.
The original <text_type> is: <original text>
Now please return the generated subsequent sentence and preceding sentence in the following format: Preceding Sentence: [The generated preceding sentence] Original Text: [The original <text_type>] Subsequent Sentence: [The generated subsequent sentence]
Prompt for Transplant
Prompt for Regeneration: The bidirectional text continuation phase generates multiple contextual scenes bearing the original text. However, these transplanted texts cannot be directly used as augmented data. Firstly, they all contain the same original text, which sacrifices the diversity of the augmented data. Secondly, these generated texts exhibit significant differences from the original text in various aspects, such as text length and format, making their differences from the original text too obvious. This is not ideal for data augmentation.
Therefore, we introduce a regeneration process, which uses LLMs to regenerate new text variants that seamlessly integrate into the expanded context. Specifically, we provide LLMs with the crafted preceding and subsequent contexts, along with the original text, and ask it to generate a new text. We implement this process using a prompt. The specific prompt template is as follows. In the prompt, we also specify some requirements for the new text generation: first, the new text should connect smoothly with the surrounding context, ensuring semantic consistency and logical flow; second, the new text should align with the original text in length, format, and linguistic style; and finally, the new text must not be a simple modification or direct copy of the original text. Instead, it should introduce novel elements, enhancing the richness and variation of the content. The prompt also explicitly instructs the LLM to adhere strictly to the original text’s label category (denoted as “<label_type>”).
1. Preceding Sentence: <preceding sentence> 2. Original <text_type>: <original text> 3. Subsequent Sentence: <subsequent sentence>
You are an expert in text data augmentation. Your task is to generate a new <text_type> that can replace the original <text_type> while meeting these requirements: 1. Fits naturally between the preceding sentence and subsequent sentence, maintaining logical flow and coherence. 2. Similar in text length, format (sentence pair, etc.), and language style to the original <text_type>. 3. Similar ’<label_type>’ to the original <text_type>, which is ’<original label>’. 4. The new <text_type> should not simply repeat the original <text_type>.
Now please return the generated new <text_type> as ’Middle Sentence’ in the following format: Preceding Sentence: [The provided preceding sentence] Middle Sentence: [The generated new <text_type>] Subsequent Sentence: [The provided subsequent sentence]
Prompt for Regeneration
Appendix B Intrinsic Evaluation
The intrinsic evaluation results based on GPT-3.5-Turbo and GPT-4o are provided in Table LABEL:tab:IntrinsicEvaluation_GPT3.5 and Table LABEL:tab:IntrinsicEvaluation_GPT4o, respectively. These results closely align with those obtained based on DeepSeek-V3 (Table 3.2), demonstrating that LMTransplant consistently generates high-quality augmentations across different LLM architectures, highlighting its robustness and effectiveness.
We also explore alternative ways of computing Distinct-N and Semantic Variability to further validate the high quality of samples generated by LMTransplant.
Distinct-N: We calculate Distinct-N across all seeds and their corresponding augmentations. However, repeated n-grams across different seeds may affect the results. To address this, we compute Distinct-N per seed and then average across all seeds. As shown in the Table LABEL:tab:Distinct-N_per_seed, this method generally yields higher scores for all augmentation methods, as it avoids counting repeated n-grams across different seeds. However, since the seeds themselves already exhibit low redundancy in n-grams (with Original’s Distinct-N close to 1.0), the overall improvement is limited. This indicates that when the seeds already have low n-gram redundancy, computing Distinct-N per seed or globally makes little difference. Additionally, this method is not applicable to GPT3Mix, as each generated text corresponds to multiple seeds, so GPT3Mix was excluded from this comparison.
Text: {text} Candidate label list: {self.label_enum_str(label_set)}
Please return ONLY the selected label name without explanations, punctuation or additional text.
Prompt for Label Prediction
The results are shown in Table LABEL:tab:LabelConsistency. We notice that LMTransplant effectively maintains label consistency in augmented texts. This demonstrates that LMTransplant can produce diverse and creative samples while preserving the core attributes of the original text, thereby maintaining the integrity and usefulness of the augmented data.
Appendix E Generate IOB Label for NER Task
For named entity recognition (NER) tasks, the IOB label sequence of each augmented text typically differs from that of the original, as the positions of named entities are likely to change during augmentation. Consequently, it is necessary to regenerate label sequences for the augmented texts. To achieve this, we employ a prompt-based approach: we provide the original text and its corresponding label sequence to LLMs, and then prompt LLMs to generate the appropriate IOB label sequence for the augmented text. The prompt template is shown below:
This task includes only the following four types of entities: persons (PER), organizations (ORG), locations (LOC), and miscellaneous names (MISC).
Use the following BIO labels and their corresponding label IDs: {’O’: 0, ’B-ORG’: 1, ’B-MISC’: 2, ’B-PER’: 3, ’I-PER’: 4, ’B-LOC’: 5, ’I-ORG’: 6, ’I-MISC’: 7, ’I-LOC’: 8}.
Please output the result strictly in the following format (only include these four lines): sentence: original sentence entities: [’token1’, ’token2’, ..., ’tokenN’] labels: [BIO_label1, BIO_label2, ..., BIO_labelN] IDs: [label_id1, label_id2, ..., label_idN]
Here is an example: sentence: {example_sentence} entities: {example_entities} labels: {example_labels} IDs: {example_ids}
Now, please perform named entity recognition and annotation for the following sentence: {input_text}
Return only the result. Do not include any explanation or additional content.
Prompt for IOB Label Generation

### v3

@@
 We choose DeepSeek-V3 as the base LLM for data augmentation.
-Results based on GPT-3.5-Turbo and GPT-4o are available in Appendix B.
+Results based on GPT-3.5-Turbo and GPT-4o are available in Appendix LABEL:sec:IntrinsicEvaluation.
 3.3 Metrics
@@
 Prompt for Regeneration
-Appendix B Intrinsic Evaluation
-The intrinsic evaluation results based on GPT-3.5-Turbo and GPT-4o are provided in Table LABEL:tab:IntrinsicEvaluation_GPT3.5 and Table LABEL:tab:IntrinsicEvaluation_GPT4o, respectively.
-These results closely align with those obtained based on DeepSeek-V3 (Table 3.2), demonstrating that LMTransplant consistently generates high-quality augmentations across different LLM architectures, highlighting its robustness and effectiveness.
-We also explore alternative ways of computing Distinct-N and Semantic Variability to further validate the high quality of samples generated by LMTransplant.
-Distinct-N: We calculate Distinct-N across all seeds and their corresponding augmentations.
-However, repeated n-grams across different seeds may affect the results.
-To address this, we compute Distinct-N per seed and then average across all seeds.
-As shown in the Table LABEL:tab:Distinct-N_per_seed, this method generally yields higher scores for all augmentation methods, as it avoids counting repeated n-grams across different seeds.
-However, since the seeds themselves already exhibit low redundancy in n-grams (with Original’s Distinct-N close to 1.0), the overall improvement is limited.
-This indicates that when the seeds already have low n-gram redundancy, computing Distinct-N per seed or globally makes little difference.
-Additionally, this method is not applicable to GPT3Mix, as each generated text corresponds to multiple seeds, so GPT3Mix was excluded from this comparison.
 Text: {text} Candidate label list: {self.label_enum_str(label_set)}